# Gerry Windiarto Mohamad Dunda Electronic and Computer Engineering

### v4

@@ IV-B Theoretical Convergence Guarantee
Define $\kappa=\frac{L}{\mu}$. Let $\eta_{w,t}=\frac{2\beta^{2}}{\mu t}$ and $\eta_{\lambda,t}=\mathcal{O}(\frac{1}{t^{2/3}})$. Given that Assumption 1 and Assumption 2 hold, each $F_{i,S}(w,\lambda)$ is $L$-smooth, each $F_{i,S}(\cdot,\lambda)$ satisfies $\mu$-PL condition, and each $F_{i,S}(w,\cdot)$ is $\rho$-strongly concave, we have
@@ Appendix A Proof of Theorem 1
$\|\nabla R_{S}(w)\|^{2}\geq\frac{2\mu}{\beta^{2}}(R_{S}(w)-\min_{w}R_{S}(w)+\Gamma-\kappa^{2}).$ (18)
@@
$=\frac{2\mu}{\beta^{2}}(F_{S}(w,\hat{\lambda}_{S}(w))-\min_{w}F_{S}(w,\hat{\lambda}_{S}(w))+\Gamma-\kappa^{2}).$
The first inequality is obtained by Jensen inequality, and the second inequality comes from the $L$-smoothness property. Since $\nabla_{w}R_{S}(w)=\nabla_{w}F_{S}(w,\hat{\lambda}_{S}(w))$ by Lemma 1, the proof is complete. ∎
@@
$\|\nabla R_{S}(w)-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}\leq L^{2}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]$
$+\frac{2L^{2}E}{\mu}(R_{S}(w_{t})-R_{S}^{*}+\Gamma).$ (19)
@@
Using Jensen inequality and $L$-smoothness, we get
$\|\nabla R_{S}(w)-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}$
$\leq\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}[\|\nabla_{w}F_{i,S}(w_{i}^{(t,0)},\hat{\lambda}_{S}(w_{t}))-\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}]$
$\leq\sum_{i=1}^{N}p_{i}L^{2}\sum_{k=1}^{E}([\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]+[\|w^{(t,0)}-w_{i}^{(t,k)}\|^{2}].$
$\leq L^{2}E\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}+L^{2}E\sum_{i=1}^{N}p_{i}\|w^{(t,0)}-w_{i}^{*}\|^{2}$
$\leq L^{2}E\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}+\frac{2L^{2}E}{\mu}\sum_{i=1}^{N}p_{i}(F_{i,S}(w_{t},\hat{\lambda}_{S}(w_{t}))$
$-F_{i,S}(w_{i}^{*},\hat{\lambda}_{S}(w_{i}^{*})))$
$\leq L^{2}E\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}+\frac{2L^{2}E}{\mu}(R_{S}(w_{t})-R_{S}^{*}+\Gamma)$ (20)
The second inequality is from $L$-smoothness. The fourth inequality is from Appendix A (PL $\rightarrow$ QG) in [28]. ∎
@@
Rewrite the combined primal update (local update + aggregation) of the global model at round $t$ ($w_{t}$) as $w_{t+1}=w_{t}-\eta_{w,t}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$, where $E$ is the number of local iterations. Assume that Assumption 1 and the conditions from Lemma 3, 2, and 1 hold. We have
$\mathbb{E}_{t}[R_{S}(w_{t+1})-R^{*}_{S}]$ $\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t}+\frac{L^{2}E\eta_{w,t}}{\mu})(R_{S}(w_{t})-R^{*}_{S})$ (21)
$+\frac{L^{2}\eta_{w,t}}{2}\mathbb{E}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]$
$+\frac{\eta^{\prime}_{w,t}}{2}E^{2}B_{w}^{2}+\frac{\mu}{\beta^{2}}\eta^{\prime}_{w,t}\kappa^{2}$
$+(\frac{L^{2}E\eta_{w,t}}{\mu}-\frac{\mu}{\beta^{2}}\eta_{w,t})\Gamma.$
@@
For simplicity, denote $w_{t}=w^{(t,0)}_{i}=w^{(t,0)}$. We start from the smoothness of $R_{S}$.
$R_{S}(w_{t+1})-R^{*}_{S}\leq R_{S}(w_{t})-R^{*}_{S}+\langle\nabla_{w}R_{S}(w_{t}),w_{t+1}-w_{t}\rangle$
$+\frac{L+\frac{L^{2}}{\rho}}{2}\|w_{t+1}-w_{t}\|^{2}$
$\leq R_{S}(w_{t})-R^{*}_{S}-\eta_{w,t}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\right\|^{2}.$
We denote $\mathbb{E}_{t}$ as the conditional expectation over all samples $\{\xi_{i}^{(t,k)}\}$ with given $w_{t}$ and $\lambda_{t}$. Taking this conditional expectation on both sides we get,
$\mathbb{E}_{t}[R_{S}(w_{t+1})-R^{*}_{S}]\leq R_{S}(w_{t})-R^{*}_{S}$
$-\eta_{w,t}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\mathbb{E}_{t}\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$
$-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})+\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})\|^{2}$
$\leq R_{S}(w_{t})-R^{*}_{S}-\eta_{w,t}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})\right\|^{2}$
@@
The last inequality is obtained by zeroing uncorrelated gradient estimators between two different clients or different local iterations, Jensen inequality, and
$\mathbb{E}_{t}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}(g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t}))\right\|^{2}\leq E^{2}B_{w}^{2}$. Further simplifying the term with $\eta_{w,t}\leq\frac{1}{(L+L^{2}/\rho)}$ and $\eta^{\prime}_{w,t}=(L+L^{2}/\rho)\eta_{w,t}^{2}$, and using Lemma 2, we obtain
$\mathbb{E}_{t}[R_{S}(w_{t+1})-R^{*}_{S}]=R_{S}(w_{t})-R^{*}_{S}$
$+\frac{\eta_{w,t}}{2}\|\nabla_{w}R_{S}(w)-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\|^{2}$
$-\frac{\eta_{w,t}}{2}\|\nabla_{w}R_{S}(w)\|^{2}\ +\frac{\eta_{w,t}}{2}E^{2}B_{w}^{2}$
$\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t})(R_{S}(w_{t})-R^{*}_{S})+\frac{\eta_{w,t}}{2}\|\nabla_{w}R_{S}(w_{t})$
$-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\|^{2}$
$+\frac{\eta^{\prime}_{w,t}}{2}E^{2}B_{w}^{2}-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\mu}{\beta^{2}}\eta^{\prime}_{w,t}\kappa^{2}$
$\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t}+\frac{L^{2}E\eta_{w,t}}{\mu})(R_{S}(w_{t})-R^{*}_{S})+\frac{L^{2}\eta_{w,t}}{2}\mathbb{E}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]$
$+\frac{\eta^{\prime}_{w,t}}{2}E^{2}B_{w}^{2}+\frac{\mu}{\beta^{2}}\eta^{\prime}_{w,t}\kappa^{2}+(\frac{L^{2}E\eta_{w,t}}{\mu}-\frac{\mu}{\beta^{2}}\eta_{w,t})\Gamma.$
The second inequality is from Lemma 3. ∎
@@
Rewrite the combined dual update as $\lambda_{t+1}=\lambda_{t}+\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}g_{\lambda}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$. Assuming the conditions from Lemma 1 and Assumption 1 hold, we have
$\mathbb{E}_{t}\|\lambda_{t+1}-\hat{\lambda}_{S}(w_{t+1})\|^{2}\leq(1+\frac{1}{\epsilon})\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}(E^{2}B_{w}^{2}$
$+2L^{2}\mathbb{E}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}])$
$+(1+\frac{1}{\epsilon})(\frac{4L^{2}\eta_{w,t}}{\rho^{2}}+\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu})(R_{S}(w_{t})-R_{S}(w^{*}))$
$+(1+\frac{1}{\epsilon})\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu}\Gamma$
$+(1+\epsilon)(1-\rho\eta_{\lambda,t})\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}+(1+\epsilon)\eta_{\lambda,t}^{2}B_{\lambda}^{2}$ (23)
@@
$\|\lambda_{t+1}-\hat{\lambda}_{S}(w_{t+1})\|^{2}$ $\leq(1+\epsilon)\|\lambda_{t+1}-\hat{\lambda}_{S}(w_{t})\|^{2}$ (24)
$+(1+\frac{1}{\epsilon})\|\hat{\lambda}_{S}(w_{t})-\hat{\lambda}_{S}(w_{t+1})\|^{2}.$
For the second term, using the fact that $\hat{\lambda}_{S}(w)$ is $\frac{L}{\rho}$-Lipschitz (Lemma 1) and applying the expectation to get
$\mathbb{E}_{t}[\|\hat{\lambda}_{S}(w_{t+1})-\hat{\lambda}_{S}(w_{t})\|^{2}]\leq\frac{L^{2}}{\rho^{2}}\mathbb{E}_{t}[\|w_{t+1}-w_{t}\|^{2}]$
$=\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}_{t}[\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\|^{2}]$
$=\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}_{t}\bigl{\|}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}(g_{w}(w^{(t,k)}_{i}-\nabla_{w}F_{i,S}(w^{(t,k)}_{i},\lambda_{t}))$
@@
$\leq\frac{2L^{2}\eta_{w,t}^{2}}{\rho^{2}}\left\|\nabla R_{S}(w_{t})-\sum_{i=1}^{N}\frac{p_{i}}{E}\sum_{k=1}^{E}\nabla F_{i}(w^{(t,k)}_{i},\lambda_{t})\right\|^{2}$
$+\frac{2L^{2}\eta_{w,t}^{2}}{\rho^{2}}\|\nabla R_{S}(w_{t})\|^{2}+\frac{L^{2}\eta_{w,t}^{2}EB_{w}^{2}}{\rho^{2}}$
$\leq\frac{2L^{4}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]+\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu}(R_{S}(w_{t})-R_{S}^{*}-\Gamma)$
$+\frac{2L^{2}\eta_{w,t}^{2}}{\rho^{2}}\|\nabla R_{S}(w_{t})\|^{2}+\frac{L^{2}\eta_{w,t}^{2}EB_{w}^{2}}{\rho^{2}}$
$\leq\frac{2L^{4}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}[\||\hat{\lambda}_{S}(w_{t})-\lambda_{t}\|^{2}]+\frac{4L^{4}E\eta_{w,t}^{2}}{\rho^{2}\mu}(R_{S}(w_{t})-R_{S}^{*}-\Gamma)$
$+\frac{L^{2}\eta_{w,t}^{2}B_{w}^{2}}{\rho^{2}}+\frac{4L^{2}\eta_{w,t}}{\rho^{2}}(R_{S}(w_{t})-R_{S}(w^{*})).$ (25)
The first inequality is obtained from Assumption 1, the second inequality is obtained from triangle inequality and independence, and the third inequality is obtained from Lemma 3.
For the first term, we have
$\mathbb{E}_{t}[\|\lambda_{t+1}-\hat{\lambda}_{S}(w_{t})\|^{2}]$
$\leq\mathbb{E}_{t}[\|\lambda_{t}+\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}g_{\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\hat{\lambda}_{S}(w_{t})\|^{2}]$
$\leq\|\lambda_{t}-\hat{\lambda}_{S}(w_{t})\|^{2}+2\eta_{\lambda,t}\mathbb{E}_{t}[\langle\lambda_{t}-\hat{\lambda}_{S}(w_{t}),$
$\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}g_{\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\rangle]$
$+\eta_{\lambda,t}^{2}\mathbb{E}_{t}[\|\sum_{i=1}^{N}p_{i}(g_{\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\nabla_{\lambda}F_{i,S}(w_{i}^{(t,E)},\lambda_{t}))$

### v5

@@ IV-B Theoretical Convergence Guarantee
Assumption 3 (Bounded Gradient).
The gradient of the local risk functions are bounded in norm by $G$,
$\|\nabla_{w}F_{i,S}(w,\lambda)\|\leq G.$ (17)
@@
Define $\kappa=\frac{L}{\mu}$. Let $\eta_{w,t}=\frac{2\beta^{2}}{\mu t}$ and $\eta_{\lambda,t}=\mathcal{O}(\frac{1}{t^{2/3}})$. Given that Assumption 1, Assumption 2, and Assumption 3 hold, each $F_{i,S}(w,\lambda)$ is $L$-smooth, each $F_{i,S}(\cdot,\lambda)$ satisfies $\mu$-PL condition, and each $F_{i,S}(w,\cdot)$ is $\rho$-strongly concave, we have
@@ Appendix A Proof of Theorem 1
$\|\nabla R_{S}(w)\|^{2}\geq\frac{2\mu}{\beta^{2}}(R_{S}(w)-\min_{w}R_{S}(w)+\Gamma)-\frac{\kappa^{2}}{\beta^{2}}.$ (19)
@@
$=\frac{2\mu}{\beta^{2}}(F_{S}(w,\hat{\lambda}_{S}(w))-\min_{w}F_{S}(w,\hat{\lambda}_{S}(w))+\Gamma)-\frac{\kappa^{2}}{\beta^{2}}.$
The first inequality is obtained by Jensen inequality, and the second inequality comes from the $L$-smoothness property. Since $\nabla_{w}R_{S}(w)=\nabla_{w}F_{S}(w,\hat{\lambda}_{S}(w))$ by Lemma 1, the proof is complete.
@@
$\mathbb{E}\|\nabla R_{S}(w_{t})-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}\leq EL^{2}\mathbb{E}[\||\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}]$
$+\eta_{w,t}^{2}L^{2}\frac{E(E+1)}{2}(B_{w}^{2}+G^{2}).$ (20)
@@
We expand the definition of $R_{S}(w_{t})$ and apply Jensen’s inequality to get
$\mathbb{E}\|\nabla R_{S}(w_{t})-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}$
$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}[\|\nabla_{w}F_{i,S}(w_{i}^{(t,0)},\lambda^{*}_{S}(w_{t}))$
$-\nabla_{w}F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\|^{2}]$
$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}L^{2}\sum_{k=1}^{E}([\||\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}]+[\|w^{(t,0)}-w_{i}^{(t,k)}\|^{2}]).$
$=\mathbb{E}L^{2}E[\|\lambda^{*}(w_{t})-\lambda_{t}\|^{2}]$
$+\eta_{w,t}^{2}L^{2}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\mathbb{E}\|\sum_{k^{\prime}=1}^{k}g_{i,w}(w_{i}^{(t,k^{\prime})},\lambda_{t};\xi_{i}^{(t,k^{\prime})})\|^{2}$
$\leq L^{2}E\mathbb{E}\|\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}$
$+\eta_{w,t}^{2}L^{2}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\sum_{k^{\prime}=1}^{k}\mathbb{E}\|g_{i,w}(w_{i}^{(t,k^{\prime})},\lambda_{t};\xi_{i}^{(t,k^{\prime})})\|^{2}$
$\leq L^{2}E\mathbb{E}\|\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}+\eta_{w,t}^{2}L^{2}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\sum_{k^{\prime}=1}^{k}(B_{w}^{2}+G^{2})$
$\leq L^{2}E\mathbb{E}[\||\hat{\lambda}^{*}_{S}(w_{t})-\lambda_{t}\|^{2}]+\eta_{w,t}^{2}L^{2}\frac{E(E+1)}{2}(B_{w}^{2}+G^{2}).$ (21)
The second inequality follows from the fact that the function is $L$-smooth. We derive the first equality by considering the $k$-th local update on the $i$-th client. The third inequality is another application of Jensen’s inequality. The fourth inequality is a consequence of Assumption 1, which bounds the variance of the gradient estimators and Assumption 3. ∎
@@
Rewrite the combined primal update (local update + aggregation) of the global model at round $t$ ($w_{t}$) as $w_{t+1}=w_{t}-\eta_{w,t}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{i,w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$, where $E$ is the number of local iterations. Assume that Assumption 1 and the conditions from Lemma 3, 2, and 1 hold. We have
$\mathbb{E}[R_{S}(w_{t+1})-R^{*}_{S}]$ $\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t})(\mathbb{E}R_{S}(w_{t})-R^{*}_{S})$
$+\frac{EL^{2}\eta_{w,t}}{2}\mathbb{E}[\||\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}]$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}+\frac{1}{2\beta^{2}}\eta_{w,t}\kappa^{2}$
$-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\eta_{w,t}^{3}L^{2}E(E+1)}{4}(B_{w}^{2}+G^{2}).$ (22)
@@
We start from the smoothness of $R_{S}$ and choose two points, $w_{t+1}$ and $w_{t}$.
$\mathbb{E}R_{S}(w_{t+1})-R^{*}_{S}$ $\leq\mathbb{E}\Big{[}R_{S}(w_{t})-R^{*}_{S}+\langle\nabla_{w}R_{S}(w_{t}),w_{t+1}-w_{t}\rangle$
$+\frac{L+\frac{L^{2}}{\rho}}{2}\|w_{t+1}-w_{t}\|^{2}\Big{]}$
$=\mathbb{E}\Big{[}R_{S}(w_{t})-R^{*}_{S}$
$-\eta_{w,t}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{i,w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{i,w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\right\|^{2}\Big{]}$
$=\mathbb{E}R_{S}(w_{t})-R^{*}_{S}-$
$\eta_{w,t}\mathbb{E}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\mathbb{E}\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{i,w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$
$-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})$
$+\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})\|^{2}$
$\leq\mathbb{E}R_{S}(w_{t})-R^{*}_{S}$
$-\eta_{w,t}\mathbb{E}\langle\nabla_{w}R_{S}(w_{t}),\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w_{i}^{(t,k)},\lambda_{t})\rangle$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t})\right\|^{2}$
@@
The first equality arises from the combined primal update. We obtain the second equality by invoking Assumption 1, which relies on the unbiasedness of the gradient estimators. The final inequality follows from the fact that we can zero out uncorrelated gradient estimators between distinct clients or local iterations, apply Jensen’s inequality, and subsequently apply the following inequality
$\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}(g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\nabla F_{i,S}(w^{(t,k)}_{i},\lambda_{t}))\right\|^{2}\leq E^{2}B_{w}^{2}$. Further simplifying the term with $\eta_{w,t}\leq\frac{1}{(L+L^{2}/\rho)}$, and using Lemma 2, we obtain
$\mathbb{E}[R_{S}(w_{t+1})-R^{*}_{S}]$ $\leq\mathbb{E}R_{S}(w_{t})-R^{*}_{S}$
$+\frac{\eta_{w,t}}{2}\mathbb{E}\|\nabla_{w}R_{S}(w_{t})-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\|^{2}$
$-\frac{\eta_{w,t}}{2}\mathbb{E}\|\nabla_{w}R_{S}(w_{t})\|^{2}\ +\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}$
$\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t})(\mathbb{E}R_{S}(w_{t})-R^{*}_{S})$
$+\frac{\eta_{w,t}}{2}\mathbb{E}\|\nabla_{w}R_{S}(w_{t})-\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\nabla F_{i,S}(w^{(t,k)},\lambda_{t})\|^{2}$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\eta_{w,t}}{2\beta^{2}}\kappa^{2}$
$\leq(1-\frac{\mu}{\beta^{2}}\eta_{w,t})(\mathbb{E}R_{S}(w_{t})-R^{*}_{S})$
$+\frac{EL^{2}\eta_{w,t}}{2}\mathbb{E}\||\lambda^{*}_{S}(w_{t})-\lambda_{t}\|^{2}$
$+\frac{(L+\frac{L^{2}}{\rho})\eta_{w,t}^{2}}{2}E^{2}B_{w}^{2}+\frac{\eta_{w,t}}{\beta^{2}}\kappa^{2}$
$-\frac{\mu}{\beta^{2}}\eta_{w,t}\Gamma+\frac{\eta_{w,t}^{3}L^{2}E(E+1)}{4}(B_{w}^{2}+G^{2}),$ (24)
where the second inequality is obtained from Lemma 2, and the third inequality is obtained from Lemma 3. ∎
@@
Rewrite the combined dual update as $\lambda_{t+1}=\lambda_{t}+\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}g_{i,\lambda}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$. Assuming the conditions from Lemma 4 and Assumption 1 hold, we have
$\mathbb{E}\|\lambda_{t+1}-\lambda^{*}_{S}(w_{t+1})\|^{2}$ $\leq(1+\frac{1}{\epsilon})\frac{L^{2}\eta_{w,t}^{2}E}{\rho^{2}}(B_{w}^{2}+G^{2})$
$+(1+\epsilon)(1-\rho\eta_{\lambda,t})\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}$
$+(1+\epsilon)\eta_{\lambda,t}^{2}B_{\lambda}^{2}$ (25)
@@
$\|\lambda_{t+1}-\lambda^{*}_{S}(w_{t+1})\|^{2}$ $\leq(1+\epsilon)\|\lambda_{t+1}-\lambda^{*}_{S}(w_{t})\|^{2}$
$+(1+\frac{1}{\epsilon})\|\lambda^{*}_{S}(w_{t})-\lambda^{*}_{S}(w_{t+1})\|^{2}.$ (26)
For the second term, using the fact that $\lambda^{*}_{S}(w)$ is $\frac{L}{\rho}$-Lipschitz (Lemma 1) and applying the expectation to get
$\mathbb{E}[\|\lambda^{*}_{S}(w_{t+1})-\lambda^{*}_{S}(w_{t})\|^{2}]\leq\frac{L^{2}}{\rho^{2}}\mathbb{E}[\|w_{t+1}-w_{t}\|^{2}]$
$=\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}[\|\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\|^{2}]$
$=\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}\mathbb{E}\bigl{\|}\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}g_{w,i}(w^{(t,k)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\nabla_{w}F_{i,S}(w^{(t,k)}_{i},\lambda_{t}))$
@@
$\leq\frac{L^{2}\eta_{w,t}^{2}}{\rho^{2}}(EB_{w}^{2}+\sum_{i=1}^{N}p_{i}\sum_{k=1}^{E}\bigl{\|}\nabla_{w}F_{i,S}(w^{(t,k)}_{i},\lambda_{t})\|^{2})$
$\leq\frac{L^{2}\eta_{w,t}^{2}E}{\rho^{2}}(B_{w}^{2}+G^{2}).$
The first equality stems from the combined primal update. We derive the second inequality from Assumption 1. The third inequality is a consequence of Jensen’s inequality, while the fourth inequality follows from Assumption 3.
For the first term, applying the combined dual update, we get
$\mathbb{E}[\|\lambda_{t+1}-\lambda^{*}_{S}(w_{t})\|^{2}]\leq\mathbb{E}[\|\lambda_{t}+\eta_{\lambda,t}\sum_{i=1}^{N}p_{i}g_{i,\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})$
$-\lambda^{*}_{S}(w_{t})\|^{2}]$
$\leq\mathbb{E}\|\lambda_{t}-\lambda^{*}_{S}(w_{t})\|^{2}+2\eta_{\lambda,t}\mathbb{E}[\langle\lambda_{t}-\lambda^{*}_{S}(w_{t}),$
$\sum_{i=1}^{N}p_{i}g_{i,\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})\rangle]$
$+\eta_{\lambda,t}^{2}\mathbb{E}[\|\sum_{i=1}^{N}p_{i}(g_{i,\lambda}(w^{(t,E)}_{i},\lambda_{t};\xi_{i}^{(t,k)})-\nabla_{\lambda}F_{i,S}(w_{i}^{(t,E)},\lambda_{t}))$